OpenStudy (anonymous):

Which expressions show how 5 • 37 can be rewritten using the distributive property? A. 5 • 30 + 7 B. 5 • 30 + 5 • 7 C. 5 • 40 – 5 • 7 D. 5 • 40 – 5 • 3 im suppose to choose multiple need help

OpenStudy (ccswims):

Are there any parenthesis?

OpenStudy (dmndlife24):

A.

OpenStudy (dmndlife24):

If you put parenthesis around the "30+7," but I'm thinking that's what you meant?

OpenStudy (anonymous):

yes thank you

OpenStudy (dmndlife24):

No problem

OpenStudy (anonymous):

Surely it's B? 5*37 = 5*(30+7) 5*(30+7)=5*30+5*7 (by the Distributive Property, as required)

OpenStudy (anonymous):

Oh, I've just seen there are multiple answers. D is also true as 37=40-3, so 5*37=5*(40-3)=5*40-5*3

OpenStudy (dmndlife24):

It's the distributive property bruhh